OpenStudy (anonymous):

given the function described by h(x)=x+8 find h(x+11)

OpenStudy (anonymous):

plug in (x+1) in for x therefore: (x+1) + 8 = x+9

OpenStudy (anonymous):

its (x+11) x+19

OpenStudy (anonymous):

so the answer is written x+19?

OpenStudy (anonymous):

whoops sorry. It looked like 1 to me so yes you are correct

OpenStudy (anonymous):

you have to change every x in function h(x) into x+11. h(x+11)= (x+11) +8 =x+19 :)